An alien spaceship moving at 91% the speed of light measure the distance between two stars to be 3.6 light years. What is the proper distance between the stars ?
1
Expert's answer
2019-04-08T09:44:55-0400
"L_0=\frac{L}{\sqrt{1-(\frac{v}{c})^2}}"

"L_0=\frac{3.6}{\sqrt{1-0.91^2}}=8.7"

The proper distance between the stars is 8.7 light years.